# Reciprocal Reward Influence Encourages Cooperation From Self-Interested Agents

### v1

@@ Abstract
Emergent cooperation among self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents. Instead, naïve reinforcement learning algorithms typically converge to Pareto-dominated outcomes in even the simplest of social dilemmas. An emerging class of opponent-shaping methods have demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents. However, they rely on higher-order derivatives through the predicted learning step of other agents or learning meta-game dynamics, which in turn rely on stringent assumptions over opponent learning rules or exponential sample complexity, respectively. To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of an opponent’s actions on their returns. This approach effectively seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without attempting to directly shape policy updates. We show that Reciprocators can be used to promote cooperation in a variety of temporally extended social dilemmas during simultaneous learning.
@@ 1 Introduction
Many species exhibit cooperative behaviors of remarkable variety and complexity. Even among prosocial animals, however, humans are especially notable for their ability to cooperate with unrelated individuals and maintain that cooperation even in highly adversarial environments (Melis & Semmann, 2010). These qualities are often credited with the development of human technology, culture, and advanced cognition (Burkart et al., 2014). As artificially intelligent (AI) agents are integrated into human society, it is increasingly important to ensure that they share these same prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
@@
However, state-of-the-art reinforcement learning (RL) methods are typically designed for the single agent setting and are ill-suited for the nonstationarities introduced by multiple agents learning simultaneously. Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce globally suboptimal outcomes. This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tensions between collective and individual return (Leibo et al., 2017). SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for fully-cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018). Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs, including mutual confession in the iterated Prisoners’ Dilemma (IPD) and an expected reward of 0 in Coins (Foerster et al., 2018).
@@
In this work, we use an intrinsic reward to encourage agents called Reciprocators to reciprocate the influence that others exert on their returns, with the goal of inducing mutually beneficial behavior from other agents. We first define a notion of influence that quantifies the effect that one agent’s action has on another’s expected return. Then, given a pair of agents, a Reciprocator $rc$ and another agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we call the influence balance that agent $i$ owes $rc$. At any time step, the Reciprocator receives an intrinsic reciprocal reward proportional to the product of its influence balance against agent $i$ and the influence of its selected action on agent $i$’s return. This incentivizes the Reciprocator to take reciprocating actions whose influence matches the sign of its current influence balance. For example, if agent $i$ has a net positive influence balance with $rc$, then $rc$’s intrinsic reciprocal reward will be positive when taking actions that positively affect agent $i$’s return.
@@
The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation by influencing others’ expected returns without modifying the reward structure of the environment, requiring higher-order derivatives, or learning meta-game dynamics. (2) We show that agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas without needing training schedules to control opponent selection. (3) Agents equipped with our intrinsic reward are able to induce cooperative behavior from purely self-interested naïve learners and several higher-order baselines.
@@ 2 Related Work
In order to improve convergence towards Pareto-optimal solutions among independently learning agents, previous work has modified agents’ reward structures to explicitly consider either group or per-capita return in order to promote cooperative behavior, e.g., via inequity aversion (Hughes et al., 2018), “empathic” harm reduction (Bussmann et al., 2019), or “altruistic” gradient adjustments (Li et al., 2024). While agents that abide by such restrictions may seek to reduce their harmful externalities, they have no way of regulating the behavior of other, less magnanimous agents, including purely self-interested “exploiters” in the worst case (Agapiou et al., 2023). Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong assumptions over the other agents’ altruistic tendencies.
@@
Another avenue towards a more robust form of cooperation relies on the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010). Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of sequential social dilemmas (SSDs). However, these methods require extensions of the original action space that allow agents to directly modify other agents’ rewards. On the other hand, we instead propose to quantify the influence of each naturally available action $a\in A$ in a given state $s$ on another agent’s expected return $R$, and use this influence as the medium for reciprocation. As we will show, such reciprocal reward and punishment can be seen as a form of “opponent shaping” that seeks to manipulate the $Q$-value of given state-action pair and consequently the likelihood of that action in future policies.
@@
Considering future policies points to a key issue with the canonical reinforcement learning (RL) framework, in that it optimizes only for the expected return within a single episode of the environment. However, taking actions that seek to optimize the long-term behavior of the other agents in the environment will not receive any immediate feedback within an episode, and must wait for one, or possibly many, learning steps. Opponent shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020, LIO), through opponents’ gradient updates (Foerster et al., 2018, LOLA), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022, MFOS). While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity to learn the dynamics of meta-games, respectively.
@@
In particular, LOLA and subsequent works such as Proximal Learning with Opponent-Learning Awareness (Zhao et al., 2022, POLA) and Consistent Learning with Opponent-Learning Awareness (Willi et al., 2022, COLA) differentiate through the opponent’s learning update but have only been demonstrated to work with vanilla gradient descent and fixed learning rates. However, modeling the learning rule of each opponent becomes increasingly implausible in the face of modern RL algorithms, which uses adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016). Indeed, LOLA with opponent modeling (LOLA-OM) showed significantly worse results than agents with direct access to their opponent’s learning rule. Another class of opponent-shaping algorithms such as MFOS uses meta-learning approaches, which must repeatedly test new meta-policies in full training runs known as “meta-episodes” in order to learn how to exploit opponent learning dynamics. This requires the ability to freely perform rollouts in the environment against opponents whose policies can be reset to initialization at will, an unrealistic assumption in environments with partially adversarial motivations.
@@
Our work is conceptually similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners. However, LOQA still differentiates through a model of the opponent’s policy and optimizes according to the joint advantage function $A(s_{t},a_{1},a_{2})$ computed with respect to the state-value function $V(s_{t})$. On the other hand, we use a counterfactual baseline that marginalizes out only the opponent’s action to perform agent-specific credit assignment and use an intrinsic reward instead of gradients over opponent policies to encourage agents to influence the opponent’s $Q$-values in the correct direction. Most importantly, LOQA focuses on the problem of learning a general end-policy that performs well against a variety of other agents at evaluation and has therefore only demonstrated the ability to shape other LOQA opponents in a controlled self-play scenario, while our method focuses on shaping the policies of diverse agents over multiple episodes of simultaneous learning.
@@ 3 Preliminaries
We construct a series of sequential social dilemmas which can each be described as a stochastic game $G$, defined by a tuple $G=\langle S,A,P,r,n,\gamma\rangle$. In $G$, $n$ agents, indexed by $i\in\{1,\ldots,n\}$, observe the state of the environment $s\in S$ and simultaneously choose actions $a^{i}\in A$ to form a joint action $\boldsymbol{a}\in\boldsymbol{A}\equiv A^{n}$. The environment then undergoes a change according to the state transition function $P(s^{\prime}|s,\boldsymbol{a}):S\times\boldsymbol{A}\times S\rightarrow[0,1]$. Each agent receives an individual reward $r^{i}=r^{i}(s,\boldsymbol{a})$, and future rewards are discounted at each time step by a discount factor $\gamma\in[0,1]$. We consider the fully observable setting where agents observe the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent. Each agent $i$ conditions a stochastic recurrent policy $\pi^{i}(a^{i}|\tau^{i})$ on its action-observation history, which is denoted as $\tau^{i}\in T\equiv(S\times\boldsymbol{A})^{*}$.
@@ 4 Reciprocal Reward Influence
Wang et al. (2020) proposed an intrinsic reward for exploration called Value of Interaction (VoI). VoI measures how much agent $i$’s actions affect agent $j$’s expected return by comparing the state-action value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ against a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$, where the $\boldsymbol{\pi}$ superscript indicates the dependence of this function on the parameters of all agents’ policies. This baseline marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return. This difference of $Q$-values can be decomposed into an immediate influence $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence computed over changes in state transition probabilities [Equation 2]. Although the original VoI was formulated as a regularizer using an expectation over trajectories $\tau$, we modify it to compute the one-step Value Influence (VI) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to compute the influence of individual actions
@@
$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\boldsymbol{\pi}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}).$ (2)
This definition of influence relies on the notion of a counterfactual baseline to assign credit to a particular agent’s action while holding all other agents’ actions constant. This counterfactual baseline is so named because it estimates the counterfactual expected return if the agent’s true action is replaced with a “default” action, which is computed by marginalizing out the agent’s action to get the expected on-policy return. This baseline return is given by
$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
In practice, we approximate this by regressing towards the Q-value while masking out $a_{i}$ from the joint state-action input. The concept of marginalizing out an agent’s action as a baseline to quantify influence has also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $i$’s expected reward that can be achieved by agent $j$’s action by taking the minimum over $a^{j}\in A$ in Equation 1. Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus, which is a special case of $VI_{i|j}$ where $i=j$ (Equation 4).
$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)

### v2

@@ Abstract
Cooperation between self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents. Instead, naïve reinforcement learning algorithms typically converge to Pareto-dominated outcomes in even the simplest of social dilemmas. An emerging literature on opponent shaping has demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents. However, such methods differentiate through the learning step of other agents or optimize for meta-game dynamics, which rely on privileged access to opponents’ learning algorithms or exponential sample complexity, respectively. To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of opponents’ actions on their returns. This approach seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without directly differentiating through a model of their policy. We show that Reciprocators can be used to promote cooperation in temporally extended social dilemmas during simultaneous learning. Our code is available at https://github.com/johnlyzhou/reciprocator/.
@@ 1 Introduction
Many species exhibit cooperative behaviors of remarkable variety and complexity. Even among prosocial animals, however, humans are especially notable for their ability to cooperate with unrelated individuals and maintain that cooperation even in highly adversarial environments (Melis & Semmann, 2010). These qualities are often credited with the development of human technology, culture, and advanced cognition (Burkart et al., 2014). As artificially intelligent (AI) agents become more commonplace in human society, it is increasingly important to ensure that they share similar prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
@@
However, state-of-the-art reinforcement learning (RL) methods are typically designed for the single agent setting and are ill-suited for the nonstationarities introduced by multiple agents learning simultaneously. Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce collectively suboptimal outcomes. This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tradeoffs between collective and individual return (Leibo et al., 2017). SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018). Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs (Foerster et al., 2018).
@@
In this work, we propose an intrinsic reward that encourages agents called Reciprocators to learn a tit-for-tat-like strategy which reciprocates the influence that others exert on their returns. We first define value influence, a notion of influence that quantifies the effect that one agent’s action has on another’s expected return. Given a pair of agents, a Reciprocator $rc$ and another learning agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we refer to as the influence balance “owed” to $rc$ by agent $i$. At each time step, the Reciprocator receives an additional intrinsic reciprocal reward proportional to the product of its current influence balance with agent $i$ and the value influence of its action on agent $i$. This encourages the Reciprocator to take actions whose influence matches the sign of the influence balance: for example, if agent $i$ has produced a net positive influence on $rc$, then $rc$ will be rewarded for actions that have a reciprocally positive influence on agent $i$’s expected return. Mechanistically, our method seeks to manipulate the opponent’s $Q$-values by altering their expected return in particular directions depending on their actions. A reward-maximizing agent $i$ should then be incentivized to take mutually beneficial actions and avoid harmful externalities.
@@
The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation from other, simultaneously learning agents without modifying the structure of the environment, computing higher-order derivatives, or learning meta-game dynamics. (2) Agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas and are able to shape purely self-interested naïve learners into mutually beneficial behavior. (3) Reciprocators demonstrate resistance to exploitation by higher-order baselines, despite using only first-order reinforcement learning algorithms.
@@ 2 Related Work
In order to improve convergence towards Pareto-optimal solutions among independently learning agents, previous work has modified agents’ reward structures to explicitly consider either group or per-capita return in order to promote cooperative behavior, e.g., via inequity aversion (Hughes et al., 2018), “empathic” harm reduction (Bussmann et al., 2019), or “altruistic” gradient adjustments (Li et al., 2024). While agents that abide by such restrictions may seek to reduce their harmful externalities, they have no way of regulating the behavior of other, less magnanimous agents, including purely self-interested “exploiters” in the worst case (Agapiou et al., 2023). Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong guarantees over the other agents’ altruistic tendencies.
2.1 Cooperation through Influence
A more robust form of cooperation can be achieved by actively exerting influence over other agents in the environment rather than unilaterally adopting a prosocial policy, especially in SSDs where extrinsic rewards incentivize defection. Jaques et al. (2019) provides an intrinsic reward for “social influence” on other agents, which is computed as the Kullback-Leibler (KL) divergence between an opponent’s policy distribution conditioned on an influence agent’s action $a^{i}$ and a counterfactual distribution which marginalizes out that influence agent. However, as a metric-agnostic quantity, KL divergence is unable to capture key details of how the action distribution changes (Park et al., 2024). In particular, this definition of influence cannot selectively modify opponents’ behavior with respect to extrinsic task returns or state transition dynamics, and therefore has limited utility as a mechanism to incentivize cooperation.
@@ 2.1 Cooperation through Influence
Other works seek to directly influence opponent returns via the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010). Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of SSDs. However, these methods require extensions of the original action space that allow agents to directly modify other agents’ rewards. On the other hand, we instead propose to quantify the influence of each naturally available action $a\in A$ in a given state $s$ on another agent’s expected return $R$, and use this influence as the medium for reciprocation. As we later show, selectively rewarding or punishing an opponent’s actions can be seen as a form of opponent shaping which seeks to manipulate the $Q$-value of given state-action pair in order to influence the likelihood of that action in future policies.
2.2 Opponent Shaping
@@
Considering future policies points to a key issue with the canonical reinforcement learning (RL) framework, in that it optimizes only for the expected return within a single episode of the environment. However, taking actions that seek to optimize the long-term behavior of the other agents in the environment will not receive any immediate feedback within an episode, and must wait for one, or possibly many, learning steps. Opponent-shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020), through opponents’ gradient updates (Foerster et al., 2018; Zhao et al., 2022; Willi et al., 2022), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022). While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity (Fung et al., 2023) to learn the dynamics of meta-games, respectively.
@@
In particular, the Learning with Opponent-Learning Awareness (Foerster et al., 2018, LOLA) class of approaches differentiate through the opponent’s learning update but have only been demonstrated to work with full access to either the opponent’s policy gradients and Hessians or their policy parameters. When using gradient approximations and modeling the opponent’s policy, LOLA with opponent modeling (LOLA-OM) showed significantly worse results, even against vanilla policy gradients. Modeling efforts become even more implausible when faced with modern RL techniques, which employ adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016) to improve learning. Model-Free Opponent Shaping (Lu et al., 2022, MFOS) was introduced as a more general meta-learning approach that requires neither privileged access to nor inconsistent assumptions on opponent learning rules. However, MFOS’s meta-policies are trained across multiple training runs repeated in sequence, each of which are treated as a single “meta-episode,” in order to learn how to exploit opponent learning dynamics. This requires the ability to freely perform rollouts in the environment against opponents whose policies can be repeatedly reset to initialization, an unrealistic assumption in environments with partially adversarial motivations.
@@
Our work is conceptually most similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners. However, LOQA still differentiates through a model of the opponent’s policy and optimizes according to the joint advantage function $A(s_{t},a_{1},a_{2})$ computed with respect to the state-value function $V(s_{t})$. On the other hand, we use a counterfactual baseline that marginalizes out only the opponent’s action to perform agent-specific credit assignment and use an intrinsic reward instead of gradients over opponent policies to encourage agents to influence the opponent’s $Q$-values in the correct direction. Most importantly, LOQA focuses on the problem of learning a general end-policy that performs well against a variety of other agents at evaluation and has therefore only demonstrated the ability to shape other LOQA opponents in a controlled self-play scenario, while our method focuses on shaping the policies of diverse agents over multiple episodes of simultaneous learning.
@@ 3 Preliminaries
We construct a series of sequential social dilemmas which can each be described as a stochastic game $G$, defined by a tuple $G=\langle S,A,P,r,n,\gamma\rangle$. In $G$, $n$ agents, indexed by $i\in\{1,\ldots,n\}$, observe the state of the environment $s\in S$ and simultaneously choose actions $a^{i}\in A$ to form a joint action $\boldsymbol{a}\in\boldsymbol{A}\equiv A^{n}$. The environment then undergoes a change according to the state transition function $P(s^{\prime}|s,\boldsymbol{a}):S\times\boldsymbol{A}\times S\rightarrow[0,1]$. Each agent receives an individual reward $r^{i}=r^{i}(s,\boldsymbol{a})$, and future rewards are discounted at each time step by a discount factor $\gamma\in[0,1]$. We consider the fully observable setting where agents have access to the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent. Each agent $i$ conditions a stochastic recurrent policy $\pi^{i}(a^{i}|\tau^{i})$ on its action-observation history, which is denoted as $\tau^{i}\in T\equiv(S\times\boldsymbol{A})^{*}$.
@@ 4 Reciprocal Reward Influence
In order to compute a general measure of value-based influence, we modify an intrinsic exploration reward proposed by Wang et al. (2020) called Value of Interaction (VoI). VoI measures how much agent $i$’s actions affect agent $j$’s expected return by computing the difference between the $Q$-value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ and a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$ which marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return [Equation 1]. Here, the bolded $\boldsymbol{\pi}$ superscript indicates the dependence of these $Q$-functions on the parameters of all agents’ policies. The VoI can be decomposed into an immediate influence term $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence term computed over changes in state transition probabilities [Equation 2]. Although the original VoI was formulated as an expectation over trajectories $\tau$ and used as a regularizer, we modify it to compute the one-step Value Influence ($VI$) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to quantify the influence of individual actions
@@
$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\pi^{i}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}),$ (2)
where $r(s,\mathbf{a}^{-i})=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[r(s,(\mathbf{a}^{-i},a^{i}))\right]$ and $p(s^{\prime}|s,\mathbf{a}^{-i})=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[p^{\pi^{i}}(s^{\prime}|s,(\mathbf{a}^{-i},a^{i}))\right]$. This definition of influence relies on the notion of a counterfactual baseline to assign credit to a particular agent’s action while holding all other agents’ actions constant. This counterfactual baseline is so named because it estimates the counterfactual expected return if the agent’s true action is replaced with a “default” action, which is computed by marginalizing out the agent’s action to get the expected on-policy return. This baseline return is given by
$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i}))\right]=\sum_{a_{i}}\pi^{i}(a_{i}|\tau^{i})Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
In practice, we approximate this by regressing towards the $Q$-value while masking out $a_{i}$ from the joint state-action input. Marginalizing out an agent’s action to quantify influence is a common paradigm in MARL, having also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $j$’s expected reward that can be achieved by agent $i$ as the minimum $VI_{i|j}^{\pi}(s,\boldsymbol{a})$ over all $a^{i}\in A$. Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus. Note that the COMA advantage function [Equation 4] can be expressed a special case of $VI_{i\mid j}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}\right)$ where $i=j$.
$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau^{i})Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)